UT 31 May 2021



Night Details

  • Sunset 19:26
  • Twilight Ends 21:00
  • Night Center 00:17
  • Twilight Begins 03:34
  • Sunrise 05:08

Night Plan

  • Starting Configuration at sunset - LUCI-Bino / LUCI-1 Authorized
  • DONE OSCO
  • DONE Bino LUCI (AO/ESM and SL) speckle-photometry 1-1.5 hours (D. Thompson) - unlikely to complete
    • /home/lbto/Service_Observing/2021A_D_E/May_Jun_Jul_2021/Bosco_2021
    • M13
    • Requires AO Support
  • BACKUP LUCI Tasks that can also be run (in case speckle-phot test has problems)
    • LUCI AO Stars (various bins)
    • Full AO Crossover test - 1 hour
      • M92 - Full AO with N375 Camera Bin 1 Ks
      • M92 - Full AO with N375 Camera Bin 3/4 Ks
    • On-Axis ESM observation Hoag's Object - 1hr
  • Handover to LBTI-Eng 22:38hr
  • Backup non-AO Eng work - in case AO (either LUCI or LBTI) can be executed
    • MODS-1 AGW Hotspot Verification - 0.5hr
    • DONE MODS Field Correction check for IT 8315 - 1-1.5 hours
      • run IDL task field_collect (same as measuring field aberrations: ComTools#Field_Data_Collection) do LDG and RDG in parallel (two IDL windows)
      • Use 6-8 stars (some fields only have 6 stars; start the list a second time, if you are under an hour)
        • Stone_G is at RA=10:17,
        • Stone_H at RA=11:03
        • Stone_J at RA=13:32,
        • and other Stone / BS fields are available if earlier or later.
    • LBC Superfocus sequences 1-2 hours
      • requires seeing < 1.2"
      • thin clouds (1-2 mags OK)
      • Grey sky
  • Poor Weather Backup (seeing >>1.2" >2 mags clouds)
    • Rothberg Merger Program (Bino LUCI)
      • /home/lbto/Service_Observing/2021A_D_E/May_Jun_Jul_2021/Rothberg_Local_Mergers
      • 1 hr each (more if conditions are REALLY pooor)
      • NGC 6598
      • NGC 4004
      • UGC 10607
      • UGC 10675
      • UGC 9829


Night Prep

Current values of guidecam_hotspot:
instr guidecam_hotspot_x guidecam_hotspot_y comment
MODS_L.cfg 254.8 199.5 y was 189.5, changed to 199.5 on 20180915
MODS_R.cfg 267.0 240.0  

Night Log

OSA: S. Allanson

ISp: Mtn Staff

Mtn Maager: M. Wagber

ISA: B. Rothberg, D. Thompson, O.kuhn (2nd half)

AO: J. Power

UT 02:20 - Opening enclosure - have thick clouds forming overhead

UT 02:34 heading over to pointing star - clouds might be thining out

UT 02:52 - still a bit bright - wating to collimate as skies darken

OSCO

UT 03:00 - starting OSCO HD 118701 - first run w/NCPA

-FWHM Guider before loop closed ~1.1"

WFSmag = 8.59 , Rmag=7.48

UT 03:15 - AO loops closed and done - but hasn't released to LUCI (IT#8205)

-aborting the LUCI script and taking the frame manually, then skipping to the next line and clicking go

(aborting script does NOT cancel the preset)

-B.Rothberg looking at IIF/TCS logs to see if anything shows up

From: /lbt/data/logs/tcs/current.events (tcs1)

Mon May 31 03:14:17.536 2021 59365.13492518578 4 AOSL aos.StartAO{FLAO}.started left StartAO {FLAO}[35]
Mon May 31 03:14:17.554 2021 59365.13492540422 3 AOSL aos.info left AO loop closed
Mon May 31 03:14:17.555 2021 59365.13492540506 4 AOSL aos.StartAO{FLAO}.complete left StartAO {FLAO}[35]
-so we know when the command was sent

in current.log:

-This is about when we manually executed the script - until this nothing else but TTF offload and H0 offload commnads in log

2021-05-31T03:20:17.022353+00:00 tcs2 LBT_AOSL: AOSL:Pause executed from IIF, handle 32
- somehow the startAO doesn't trigger anything on LUCI.

Speckle-Phot test

L1 0004-0042

UT 03:32 - authorizing for bino-LUCI

-opening the AO loop

UT 03:36 - sending preset to M13

Rmag = 11.98, WFSMag = 13.33

UT 03:47 - started

from event.log:

Mon May 31 03:46:55.400 2021 59365.15758565018 4 AOSL aos.StartAO{FLAO}.started left StartAO {FLAO}[50]
Mon May 31 03:46:55.412 2021 59365.15758579638 3 AOSL aos.info left AO loop closed
Mon May 31 03:46:55.412 2021 59365.15758579721 4 AOSL aos.StartAO{FLAO}.complete left StartAO {FLAO}[50]
- so we see the startAO but no confirmation in the log - need to find if there is a log somewhere where LUCI "sees" that StartAO

UT 03:48 - D.Thompson noticed some odd motor motions on FM4 - fieldstop on both LUCIs moved

L1 0043 FS is noticeable off

and for L1 044 and L2 006 moved back into a better position

UT 03:53 - loops opened after offset L1 0047 do not use

UT 03:54 - re-optimizing gain

UT 03:57 - loops closed - continuing

UT 04:05 - switched to sub-framing - ESM

UT 04:11 - Loops closed

UT 04:24 - some thin clouds passing over

UT 04:34 - sky frame - we happen to be pointing exactly at a cloud (should reduce the flux of the stars...)

UT 04:40 - offset failed left side, right side (no star in AO WFS) - right side seems "stuck"

-IIF is stuck - not sure what to do on right side

-had to release the synch offset on right

UT 04:46 - stoping AO - giving up on this object, will move to an ESM target

L1 0043 - 0050

L1 0005- 0014

ESM Test Hoags Objects

UT 04:49 - sending preset

-sent wrong preset - L2 was TTM too

-third time's the charm

UT 04:57 - worked, but we hit the edge of the cloud and no guidestar

UT 05:18 - still in clouds ....

UT 05:21 - Relinquising to LBTI

UT 05:22 - putting LUCIs into safe config

UT 05:23 - reconfiguring for LBTI

UT 07:14 - reconfiguring for bino-MODS

Adjust AGw hotspot on MODS1

  • Relaxed the tolerance on the maxOffsetToHotspot and maxHotspotFromProbe to their values + 2.
Current values of guidecam_hotspot:
instr guidecam_hotspot_x guidecam_hotspot_y comment
MODS_L.cfg 254.8 199.5 y was 189.5, changed to 199.5 on 20180915
MODS_R.cfg 267.0 240.0  
07:17 Steve is going to a 15th mag star for the check

gcs_wfsc_display used

Between WFS image 197 and 198, I issued the command setHotspot left -x 254.8 -y 195 (-4 mm in Y)
  • 199 looks better than 197, but seeing is better
Between WFS image 205 and 206, setHotspot left -x 254.8 -y 199.5
  • 207 looks good - this was the original
Made a big offset - setHotspot left -x 254.8 -y 189 (star seems to be on the right)
  • after 222, setHotspot left -x 254.8 -y 180 (moved to the right edge, see 225)
Moved back to setHotspot left -x 254.8 -y 195
  • after 230, setHotspot left -x 254.8 -y 210 (moved to left edge)
  • 234, setHotspot left -x 254.8 -y 215 (further to left edge)

setHotSpot left -x254.8 -y 201 looks good. Settling on that. Updating the MODS_L.cfg file.

For MODS2, all looks good with the guider_hotspot, set tolerances back.

Updated values of guidecam_hotspot are: Current values of guidecam_hotspot:
instr guidecam_hotspot_x guidecam_hotspot_y comment
MODS_L.cfg 254.8 201  
MODS_R.cfg 267.0 240.0  
08:03 Steve bounced GCS and is presetting back to this target. The seeing has been ~0.7-0.8" during this adjustment.

Align hotspot to rotator center (take data only, no changes made)

Steve sent an active preset to the same star and stopped guiding after it was collimated.

ptincrement 0,1 CA -5

gprobe 0 0

gcsclient readGuideCam left,right -e 90000

slewtohold 360 -> 0 for image

On the images, the circles are centered at:
/lbt/data/share/tcs/lbto.20210531/left_guider000001.fits 179 173
/lbt/data/share/tcs/lbto.20210531/right_guider000001.fits 185 162

When these centers are corrected for the windowing, they become:

  • sx detsec [76:447,22:379] --> 256, 196 for SX cf 254.8, 201
  • dx detsec [85,449,51,424] ---> 271, 214 for DX cf 267, 240

The differences, HS-RC are:
  xc yc xhs[pix] yhs[pix] (xc-xHS)[pix] (yc-yHS)[pix] (xc-xHS)[mm] (yc-yHS)[mm] (xc-xHS)["] (yc-yHS)["] Comments
MODS1 254 194 254.8 201 -0.8 -5 -0.051 -0.319 -0.085 -0.53 GP travel to +Y may run into limit
MODS2 269 212 267 240 2 -28 0.127 -1.783 0.212 -2.968 RC is ~3" below RC on guider image
  • 0.106"/pix and
  • 0.0637 mm/pix
  • 0.6 mm/"
DX side is pretty far off in Y so I am repeating.

Collimating on-axis
Steve stopped guiding
I sent the rotators to 360
ptincrement 0,1 CA -5
gprobe 0 0
gcsreadGuideCam left,right -e 90000
slewtohold (0->360)

/lbt/data/share/tcs/lbto.20210531/left_guider000002.fits 180 172
/lbt/data/share/tcs/lbto.20210531/right_guider000002.fits 186 159

The rotator centers are consistent but on the DX side, the guider_hotspot is offset from the rotator center in Y by ~30 pixels or 3".

Some questions:
  • After the preset, the initial guide probe positions are not 0 0 and this is explained as due to the refraction corretion. But the deltas are different for MODS1 and MODS2.
  • Could this have to do with the pointing origins that are set in PCSInstrument.conf, which are not 0 0 for MODS but are:
Instr PointingOrigin X PointingOrigin Y
MODS1 -1.90 -1.20 alignment stars nearly centered in alignment boxes
MODS2 -3.75 -1.00 typical offset (dtheta,dx,dy) to align mask are below

Statistics from MOS alignments in 2020-2021 (sig_N-1 in parens - no sigma-clipping):
Instr stat dtheta dx dy
MODS1 avg -0.17(0.08) -0.47(0.90) 1.31(1.49)
  med -0.17 -0.59 0.99
MODS2 avg -0.03(0.07) 3.59(1.53) -1.01(0.76)
  med -0.04 3.87 -1.00

Collection of data to measure field aberrations

09:00 Field Aberration Measurements

09:11 Preset to BS9177:field_collection. Seeing 0.81" on the DIMM

sxfield/BS9177_field.log
#| GStar | wfsc # |
2021531
BS9177_field
wfsc
GCS
| 0 | 273-274 |
| 11 | 275-277 |
| 0 | 278-279 |
| 16 | 280-282 |
| 0 | 283-285 |
| 13 | 286-288 |
| 0 | 289 |
| 17 | 290-292 |
| 0 | 293-294 |
| 21 | 295-297 |
| 0 | 298-300 |
| 6 | 301-303 |
| 0 | 307-306 |

#| GStar | wfsc # |
2021531
BS9177_field
wfsc
GCS
| 0 | 180-181 |
| 11 | 182-184 |
| 0 | 185-187 |
| 16 | 188-190 |
| 0 | 191 |
| 13 | 192-194 |
| 0 | 195-196 |
| 17 | 197-199 |
| 0 | 200-201 |
| 21 | 202-204 |
| 0 | 205-207 |
| 6 | 208-210 |
| 0 | 211 |
| 3 | 212-214 |

LBC Superfocs with TMS passive

9:45 Heading up to zenith and reconfiguring to the LBC's for superfoc sequences

10:07 Preset to F3 Copointing star and running do hybrid

10:19 blue reached convergence with 0.95 seeing, red reached convergence with 0.67". Olga is taking a reference to run TMS in passive. Running LBC rangebal

10:25 Running RB_rv_sf7. (PA0). Seeing 1.05" on the DIMM. Striping evident in red science images.
10:32 Running RB_rv_sf7_pa180. (PA0). Seeing 1.25" on the DIMM
10:40 Running RB_rv_sf7. (PA0). Seeing 1.31" on the DIMM.

10:46 running dofpia again. The blue images look aberrated beyond just focus. DIMM 1.22"
Red converged with 0.84", blue converged with 1.03"

10:51 Running RB_rv_sf7. (PA0). Seeing 1.14" on the DIMM.
10:57 Running RB_rv_sf7_pa180. (PA0). Seeing 1.13" on the DIMM
11:04 Running RB_rv_sf7. (PA0). Seeing 1.25" on the DIMM.

11:13 Preset to F5 CoPointing and running dofpia.

11:17 Running copointing

11:20 Starting RB_rV_sf7 on F5 at PA0. Seeing 1.23" on the DIMM.
11:27 Starting RB_rV_sf7_pa180 on F5 at PA180. Seeing 1.29" on the DIMM.

11:34 slewing for test for skyflat. B & R. PA 0 & PA 180

11:49 Trying for some Uspec and Y. Scaling is off so changing from 1 to 1.22 -> 1 to 3. Only PA 180 obtained.

-- BarryRothberg - 18 May 2021
Topic revision: r20 - 01 Jun 2021, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback